Week of 061204

Open Actions from last week:
  • Check in opm grid templates for reference to prod-grid-cern-admins (James) - In Progress
  • FTS Intevention (Gavin) - DONE

Chair: James

SMOD: Ulrich

GMOD: Yvan (Maite)


Log: Power cut to RB/CE nodes this morning during scheduled power test. Dteam rate to 0 since 5pm sunday - to investigate

New Actions:

  • CRL update system change on all nodes today ( Jan) - DONE
  • Actuator needed for CE Grid_Gris_Wrong (Thorsten) - DONE
  • Check with Lyon on SRM availability for Multi-VO tests (Harry) - DONE
  • Grid service contract type check (and change) (James) - In Progress


  • FTS Intervention will happen today now that lxservb02 is back


Log: Castor problem on castorpublic cluster - no dteam transfers being scheduled since 11pm last night

New Actions:


  • IN2P3 will do hardware upgrade tomorrow. Then ready for multi-vo tests. FZK have said maybe they can do something earlier (pre 1.7 upgrade on 12th) - will get back this morning to us.


Log: Dteam on castor ok again



  • Castor problems diagnosed - DLF problems for dteam. DB lock for Alice.


Log: Lots of Errors in FTS pages for GSSDATLAS.


  • follow up errors with GSSDATLAS (Simone/Harry)


Log: rb113 high load. CMS started transfers to Lyon, but problems with SRM at Lyon


  • up high load limit on rb113 (Yvan)


Edit | Attach | Watch | Print version | History: r5 < r4 < r3 < r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r5 - 2007-02-02 - FlaviaDonno
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    LCG All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback